Station ASCN, Butt Crater, Ascension Island

ASCN operational status:

2024.05.19

IDA engineers visited the station and replaced the T120PH with an STS-5 and installed a cell modem.

2023.08.27

The station is back online after the station operator swapped the fan motor from the shipped Newmar into the  Newmar that was overheating (the new Newmar had a faulty display).  An upgraded Newmar will still be shipped to the station.

2023.08.19

The station is down due to a faulty Newmar power supply.  A new Newmar has been shipped to the station.

2023.07.24

Data quality from the T120PH is poor after the abrupt drop / offset in counts on day 2023,205-17:09.

2022.03.03

IDA engineers are onsite to install a Trillium 360BH seismometer as the primary sensor (replacing decommissioned KS54000 and faulty STS-6).  Both station computers were replaced.

2021.12.01

The STS-6 installed at the station is faulty.  It will be returned to UCSD for testing and repairs.

2021.11.22

IDA engineers are onsite to replace the KS54000 seismometer with an STS-6,  replace the wellhead fiberglass enclosure, run fiber optic line from the wellhead to the recording room, update the Q330 firmware, and perform other station maintenance.

2021.10.20

The station operator restored communications between the wellhead and recording room.

2021.09.15

The station is down as of day 2021,258-08:43 due to a damaged cat-5 cable (chewed by rats) between the wellhead and recording room.

2021.01.27

The station operators installed new cat-5 cable in PVC conduit from the wellhead to the recording room.  The cat-5 cable had been repeatedly chewed by rats which was causing intermittent communication issues and data loss since November 30,2020.

2019.03.29

All channels from the KS-54000 are non-seismic as of day 2019,088-21:42.  The vertical channels have been non-seismic since 2012.

2017.07.12

The station was down from day 2017,187-20:47 to 2017,191-10:05 due to power issues in the equipment rack.  Data acquisition was restored after the station operator visited the station and repaired a damaged network switch cable, and power supply.

2013.09.09

IDA engineers visited the station and installed a Trillium posthole seismometer to replace the CMG-3T seismometer.

2012.11.20

The vertical channels from the KS-54000 became non-seismic on day 2012,325-23:24 for reasons yet to be determined. The vertical channels started showing a strange signal for about an hour a day beginning 2012,323 before becoming non-seismic.

2010.10.20

The station is back on-line after the station operator replaced the ISP.

2010.08.26

The station is down as of day 2010,238-13:49 due to problems with the ISP. A new station computer is being shipped to the station.

2010.06.30

The station was down from day 2010,165-09:15 to 2010,181-10:14 due to problems with the station computer. Data acquisition was restored after the station computer was rebooted and several processes were restarted.

2010.05.24

The station operator replaced the seismometer interface box on day 2010,132 (May 12). Remote calibration of instruments at the station is now possible.

2009.05.23

The station is up after an IDA engineer visited the station and upgraded it to a Quanterra Q330HR system.

2009.03.15

The station is down since about day 2009,073 due to problems with the DAS that are yet to be determined.

2009.01.12

ASCN timing issue: The timing system at ASCN began behaving strangely on 2007:320-10:51. When this occurred, clock lock was lost, so the DAS’ oscillator was no longer adjusted to the nominal data rate. After repeated attempts to resolve the problem with the assistance of the station operator, an IDA engineer traveled to the site from June 3-16, 2008. He replaced the DAS and clock but was unable to resolve the problem, which is now believed to be rooted in a grounding issue along the uplink cable. We are now planning to upgrade the data acquisition system and replace the uplink cable sometime in early 2009. Using the station’s computer whose own oscillator is disciplined by NTP, it is possible to repair the timing of the data to an accuracy of less than one minute. Because the quality of the waveform data remains high, we have elected to distribute these data with bit 7 of the data quality flag in the fixed section of the SEED data header set to indicate that the time tag is questionable. Data from 2008,177 (June 25th) are distributed in this way. Peter Davis, January 12, 2009

2008.06.12

The north-south channels from the CMG3-T seismometer are non-seismic. The sensor needs to be recentered.

2008.04.07

The station operator replaced the DAS and GPS clock on day 2008,094. However, the station is still experiencing problems with the new DAS and GPS clock.

2007.11.16

The timing for data from IDA station ASCN is unreliable as of about day 2007,320 due to problems with the GPS clock and the DAS that are being investigated.

2006.05.23

The vertical channels for the CMG3-T seismometer were restored about day 139-14:26 after the station operator recentered the masses for the CMG3-T seismometer.

2006.04.12

The vertical channels for the CMG3-T seismometer are out. The station operator is requested to recenter the masses for the CMG3-T seismometer.

2004.11.05

Station is back up after the station operator replaced the power supply, LAN extender, and GPS clock. Telemetered data acquired as of day 309-13:17(Nov. 4).

2004.09.20

Power supply equipment in transit to station.

2004.08.19

Station is down due to problem with vault power supply. No data since 8/5/2004.

2004.08.09

Telemetry still down. Station operator asked to power cycle DAS and modem, and check modem connections.

2004.07.12

Telemetry data are still not available. The station operator has been asked to power cycle the ISP computer.

2004.07.09

Station operator cycled power on modem and reported that seismic equipment appears to be operating normally.

2004.07.08

Telemetry data stop coming in since 2004184. Station operator has been asked to check the system situation, the connectivity status and power cycle the modem.

2003.12.01

There are connectivity problems to the station as of 11/28, due to problems with the station’s modem.

2003.11.24

The data acquisition system has been upgraded to a MK8 system. Additional sensors have been added including CMG-3T and FBA-23 seismometers and a Paroscientific microbarograph.

2003.09.19

Station is down due to problems with the DAS.

2003.05.16

BBZ channels back to normal as of 2003:096-13:55 (April 6, 2003).

2003.03.12

Vault power supply fixed. SP,MP channels dead, BBZ channel dead as of 2003:069-08:43 (March 10,2003).

2002.11.28

New EPROMs installed to correct DSP problem.

2002.11.25

Station is back up after replacement of power supply.

2002.11.05

Station is down due to power supply problem.

2002.04.22

Data streams from STS2 are dead. Recent tapes are difficult to read so a replacement DAT has been shipped to the station.